OpenStudy (anonymous):

If the product of all the unique positive divisors of n, a positive integer which is not a perfect cube, is n^2, then the product of all the unique positive divisors of n^2 is (A) n^3 (B) n^4 (C) n^6 (D) n^8 (E) n^9 - - i got E. is that right?

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

Can someone explain how to do this problem? Really curious as of why that is the answer.

OpenStudy (nincompoop):

?

OpenStudy (ikram002p):

weird question

OpenStudy (ikram002p):

n^2= n*n or n^2=n^2 *1

OpenStudy (ikram002p):

so hehe since n^2 * n^2 * (n^2)^2 *1=n^8 how did u got 9 ?

ganeshie8 (ganeshie8):

*

OpenStudy (ikram002p):

ok lol i forget one factor xD factors of n^2 are 1 n n^2 factors from n there multi = n^2 factors from n^2 with multi =(n^2 )^2=n^4 so we have 1*n*n^2 * n^2 *n^4 =n^9 cool ^^

ganeshie8 (ganeshie8):

nice :)

OpenStudy (nincompoop):

no one trusts me :(

OpenStudy (ikram002p):

lol why :o